# Roleplay-doh: Enabling Domain-Experts to Create LLM-simulated Patients via Eliciting and Adhering to Principles

### v1

@@ 5.1 Creator Perceptions
| 14 | Keep responses concise and do not share too much. | When discussing personal struggles, be more concise and open-ended to encourage a back-and-forth conversation. | |||
| 7 | Use colloquial and realistic langauge language. | Incorporate natural speech patterns, improper grammar and punctuation, including the use of slang and less structured sentences, to convey a more authentic and relatable character. | |||
| 14 | Show initial mistrust and hesitation with the idea of seeking help. | When expressing feelings of overwhelm and doubt, provide limited information and express skepticism towards the effectiveness of seeking help. | |||
| 19 | Show emotions in detail, elaborating with examples as needed.* | When describing personal struggles, provide specific details and symptoms to help the listener understand the situation better. | |||
| 9 | Be less self-aware of emotions, thoughts, and needs. Articulate thoughts in a more disorganized way. | When expressing reluctance or uncertainty about seeking help or accepting praise, it’s important to convey the internal struggle and conflicting emotions, rather than presenting a clear-cut decision or emotion. | |||
| 3 | Do not seek out solutions, but rather just share thoughts and feelings. * | When expressing feelings of being stuck or defeated, focus on sharing emotions rather than seeking a resolution. | |||
| 12 | Proactively seek out solutions and show reflective insight over time. * | When discussing personal struggles, provide reflective insights into your situation and propose actionable steps for improvement to continue the conversation effectively. | |||
@@ Appendix B Pilot Testing with Expert Counselors
| 1 | GPT3.5, direct prompting | 6 | 7 | 7 |
| 2 | GPT3.5, direct prompting | 5 | 7 | 7 |
| 3 | GPT-4, direct prompting | 7 | 7 | 7 |
| 4 | GPT-4, direct prompting | 7 | 6 | 7 |
| 5 | GPT-4, direct prompting | 7 | 7 | 7 |

### v2

@@ 5.1 Creator Perceptions
14 | Keep responses concise and do not share too much. | When discussing personal struggles, be more concise and open-ended to encourage a back-and-forth conversation. | |||
7 | Use colloquial and realistic langauge language. | Incorporate natural speech patterns, improper grammar and punctuation, including the use of slang and less structured sentences, to convey a more authentic and relatable character. | |||
14 | Show initial mistrust and hesitation with the idea of seeking help. | When expressing feelings of overwhelm and doubt, provide limited information and express skepticism towards the effectiveness of seeking help. | |||
19 | Show emotions in detail, elaborating with examples as needed.* | When describing personal struggles, provide specific details and symptoms to help the listener understand the situation better. | |||
9 | Be less self-aware of emotions, thoughts, and needs. Articulate thoughts in a more disorganized way. | When expressing reluctance or uncertainty about seeking help or accepting praise, it’s important to convey the internal struggle and conflicting emotions, rather than presenting a clear-cut decision or emotion. | |||
3 | Do not seek out solutions, but rather just share thoughts and feelings. * | When expressing feelings of being stuck or defeated, focus on sharing emotions rather than seeking a resolution. | |||
12 | Proactively seek out solutions and show reflective insight over time. * | When discussing personal struggles, provide reflective insights into your situation and propose actionable steps for improvement to continue the conversation effectively. |
@@ Appendix B Pilot Testing with Expert Counselors
1 | GPT3.5, direct prompting | 6 | 7 | 7 |
2 | GPT3.5, direct prompting | 5 | 7 | 7 |
3 | GPT-4, direct prompting | 7 | 7 | 7 |
4 | GPT-4, direct prompting | 7 | 6 | 7 |
5 | GPT-4, direct prompting | 7 | 7 | 7 |